# Exploring the unleaved tree of numerical semigroups up to a given genus

### v2

@@ 1 Introduction
Given a numerical semigroup $\Lambda$, consider the conductor $c(\Lambda)$, which is the first non-gap larger than the Frobenius number, the multiplicity $m(\Lambda)$, which is its first non-zero non-gap, the jump $u(\Lambda)$, which is the difference between the second and the first non-zero non-gaps, and the efficacy $r(\Lambda)$, which is the number of right generators. An encoding of a numerical semigroup will be a finite set of finite parameters uniquely defining the numerical semigroup, together with its conductor, miltiplicity, jump, efficacy and genus. We will describe a general framework for algorithms exploring the semigroup tree up to a given genus for a general encoding. This framework fits the previous RGD-algorithm [9] and the seeds algorithm [8, 6]. It just involves the notions of right generators, strong generators (defined in Section 2), pseudo-ordinary and quasi-ordinary semigroups.
Then we will present an encoding system, that, used in the previous farmework, will allow to trim all the leaves and branches of ${\mathscr{T}}$ not arriving to a given depth, providing the so-called unleaved tree of numerical semigroups of a given depth (see Figure 1). This represents a new paradigm in the exploration of the semigroup tree up to a given genus, which will lead to much more efficient algorithms than the ones used up to date. It has been already used to compute the number of semigroups of genus $76$ and $77$:
@@ 5 More unvisited nodes
[15, Corollary 10] There is exactly one semigroup of genus $\gamma$ and multiplicity $2$ and there are $\gamma-\lfloor\frac{2\gamma-1}{3}\rfloor$ semigroups of genus $g$ and multiplicity $3$

### v3

@@ 1 Introduction
Nowadays, prior to this work, the RGD-algorithm [9] is the fastest implemented algorithm capable of computing $n_{g}$ with no bounds on $g$, the seeds algorithm [8, 6] is the fastest implemented algorithm but limited to genera up to around one half of the maximum size of native integers. On its side, the Fromentin-Hivert algorithm [13] is the one that has been used to compute $n_{g}$ up to a largest genus (up to genus 75) [12].
Given a numerical semigroup $\Lambda$, consider the conductor $c(\Lambda)$, which is the first non-gap larger than the Frobenius number, the multiplicity $m(\Lambda)$, which is its first non-zero non-gap, the jump $u(\Lambda)$, which is the difference between the second and the first non-zero non-gaps, and the efficacy $r(\Lambda)$, which is the number of right generators. An encoding of a numerical semigroup will be a finite set of finite parameters which, together with its conductor, multiplicity, jump, efficacy and genus, uniquely define the numerical semigroup. We will describe a general framework for algorithms exploring the semigroup tree up to a given genus for a general encoding. This framework fits the previous RGD-algorithm and the seeds algorithm. It just involves the notions of right generators, strong generators (defined in Section 2), pseudo-ordinary and quasi-ordinary semigroups.
Then we will present an encoding system, that, used in the previous farmework, will allow to visit all nodes at a given depth while trimming all the leaves and branches of ${\mathscr{T}}$ not arriving to that depth, providing the so-called unleaved tree of numerical semigroups of the given depth (see Figure 1). This represents a new paradigm in the exploration of the semigroup tree up to a given genus, which will lead to much more efficient algorithms than the ones used up to date. The idea of pruning branches of the tree has been used in the context of the search of counterexamples of Wilf’s conjecture (or to the related Eliahou’s conditions) in [11, 12], in which nodes whose descendants are known to verify the conjectures are not futher explored.
Our algorithm has been already used to compute the number of semigroups of genus $76$ and $77$:
@@ 5 More unvisited nodes
[16, Corollary 10] There is exactly one semigroup of genus $\gamma$ and multiplicity $2$ and there are $\gamma-\lfloor\frac{2\gamma-1}{3}\rfloor$ semigroups of genus $g$ and multiplicity $3$
@@ 6 The algorithm ExploreUnleavedTree
7 An unleaved version of the seeds algorithm
The seeds algorithm presented in [8] and improved in terms of its implementation in [6], can also be simplified by means of visiting only the unleaved tree. The main difference with respect to the algorithm explained so for is that the shrunk semigroup of a given semigroup (as well as the seeds table) is encoded as the set of 1-bits in a long integer. Since shrinking and amplifying the bit sets in the long integer is not convenient, the shrunk semigroup is kept as just the monoid generated by the left elements of the semigroup, without dividing by the gcd of the left elements. The gcd is still computed in order to discriminate when the genus of the shrunk semigroup needs to be compared to $\gamma$.
Its implementation can be found in…
It turns out that for some genera the unleaved version of the seeds algorithm is faster than the algorithm explained in the previous sections, but not for all genera. In any case, they both perform much faster than any of the previous algorithms in the current literature.